# How Global Calibration Strengthens Multiaccuracy

### v1

@@ 2.2 Hardcore Sets of Optimal Density
A measure on $\mathcal{X}$ is a function $\mu:\mathcal{X}\rightarrow[0,1]$, which is not identically 00. It defines a probability distribution $\bar{\mu}$ by re-weighting $\mathcal{D}_{\mathcal{X}}$ by $\mu$. A hardcore measure $\mu$ is one which induces a hardcore distribution $\bar{\mu}$. Given a measure $\mu:\mathcal{X}\rightarrow[0,1]$ and a base distribution $\mathcal{D}_{\mathcal{X}}$ on $\mathcal{X}$, the density of $\mu$ in $\mathcal{D}_{\mathcal{X}}$ is given by
@@
IHCL+++++ + from multicalibration.
Recently, Casacuberta, Dwork and Vadhan [8] showed that starting from a multicalibrated predictor, one can derive a stronger and more general version of the IHCL, which they call IHCL+++++ +. Specifically, they show that given a multicalibrated partition, we can construct hardcore distributions of locally optimal density on each (large enough) piece of the partition. As expected, given that multicalibration is stronger than multiaccuracy, one can recover the original IHCL from IHCL+++++ + by “gluing” these local hardcore distributions. This glued distribution is a global $2\delta$-dense hardcore measure.
@@
We show how to construct hardcore distributions with optimal density $2\delta$ as in [8] but with the lower complexity bound of $q=O(1/(\varepsilon^{2}\delta^{2}))$ for $\mathcal{C}_{q}$ as in [46]. Indeed, we give a simple and explicit form for such a measure, starting from a predictor satisfying calibration and a weighted variant of multiaccuracy that we call weighted multiaccuracy. Both calibration and weighted multiaccuracy are implied by multicalibration, so this gives an alternate proof of the optimal density result of [8]. But crucially, the cost of obtaining both these conditions together (by many measures, including oracle calls to the weak learner) is not much more than that of multiaccuracy, which yields a considerably better bound on the parameter $q$.
To formally state our result, we define weighted multiaccuracy. We will consider weight functions $w:[0,1]\to[1,2]$, which assign weights to predictions, as is standard in the calibration literature [19, 3]. While the range $[-1,1]$ is more standard, this choice of range is appropriate for hardcore measures. The weight function that will be important for us is
@@
Along the way, our argument fixes a gap in the gluing argument used to derive the IHCL with density $2\delta$ from IHCL+++++ + in [8], see the discussion at the end of Section 5.3.666Gluing is straightforward for hardcore sets, but more subtle for hardcore measures.
@@
We explain at an intuitive level how $\mu_{\mathsf{Max}}$ improves over $\mu_{\mathsf{TTV}}$ and gets optimal density. Consider a level set $S_{v}=\{x:p(x)=v\}$ of the predictor $p$ and partition it into $S_{v}^{0}$ and $S_{v}^{1}$ based on whether $g(x)$ is 00 or $1$. $\mu_{\mathsf{TTV}}$ assigns weight $1-v$ to the points in $S_{v}^{1}$ and $v$ to points in $S_{v}^{0}$. If the predictor $p$ is calibrated for this level set, so that $\Pr[g(\mathbf{x})=1|\mathbf{x}\in S_{v}]=v$, then
@@
Crucially, the measure gives equal weight $v(1-v)$ to $S_{v}^{0}$ and $S_{v}^{1}$.
Now consider the measure $\mu_{\mathsf{Max}}$. If $v<1/2$ then $\mu_{\mathsf{Max}}$ assigns weight $1$ to points in $S_{v}^{1}$, which are in the minority, and constitute a $v$ fraction of $S_{v}$. It assigns weight $v/(1-v)$ to points in $S_{v}^{0}$, which are in the majority, and constitute a $1-v$ fraction of $S_{v}$. Overall the weight given to each of $S_{v}^{0}$ and $S_{v}^{1}$ is $v=\min(v,1-v)$. When $v>1/2$, the set $S_{v}^{1}$ has the majority of points, so $\mu_{\mathsf{Max}}$ assigns weight $1$ to points in $S_{v}^{0}$, and weighs $S_{v}^{1}$ so that both have equal weight $\min(v,1-v)$.
Why is this a good weighting strategy? Consider the majority predictor $\tilde{p}=\mathrm{1}(p\geq 1/2)$. [^7] This predictor is the best-response if $p$ is Bayes optimal. $\mu_{\mathsf{Max}}$ assigns weight $1$ to the points where $\tilde{p}$ is incorrect. The rest of the weights are assigned so that $S_{v}$ is balanced, assuming that $p$ is in fact calibrated. This clearly ensures that $\tilde{p}$ is incorrect with probability $1/2$. We now want to argue both that $\mu_{\mathsf{Max}}$ has good density and that no hypothesis in $\mathcal{C}$ can do better than $1/2$. It is intuitive that if $\mu_{\mathsf{Max}}$ is hard for $\tilde{p}$, which incorporates the distinguishing power of the class $\mathcal{C}$, then it should also be hard for $\mathcal{C}$.
@@
It remains to argue that the measure $\mu_{\mathsf{Max}}$ is hardcore. In Lemma 5.8 we present a characterization of the prediction advantage of $\mathcal{C}$ under the distribution $\bar{\mu}_{w}$ in terms of the $(w,\mathcal{C})$-weighted multiaccuracy error of $p$ under the distribution $\mathcal{D}^{g}$. This tells us that if the multiaccuracy error is $\tau$, then the prediction advantage under $\bar{\mu}_{w}$ is $\tau/\mathsf{dns}(\mu_{w})$. This proof technique is different from [8, Lemma 2.17] who show that if a function is indistinguishable from a constant function $v$ to $\mathcal{C}$ on each level set $S_{v}$, then then one can obtain a hardcore set of density $2\min(v,1-v)$ within $S_{v}$. The required indistinguishability holds for every level set by the assumption of multicalibration. Our result can be seen as a simplifcation and extension of the hardness result in [46] to weighted measures.
Note that calibrated multiaccuracy does not give the stronger IHCL+++++ + theorem proved in [8]; indeed that seems to require the full power of multicalibration. Also, the min-max based proof of Nisan as cited in [27] and the boosting-based proof of [35] give $q$ which grows as $O(\log(1/\delta))$, but they do not get optimal density. The $1/\varepsilon^{2}$ dependence on $\varepsilon$ was recently showed to be necessary by [6].
@@ 2.3 Multiaccuracy and Restricted Weak Agnostic Learning
  • The Blue line with slope $1$ follows from the boosting results of [30, 13].
@@ 3.1 Agnostic Learning
The transformation $\mathbf{y}\to 2\mathbf{y}-1$ maps $\{0,1\}$ to {±$\{\pm 1\}$. Hence the correlation measures the correlation between $c$ and {±$\{\pm 1\}$ labels associated with $\mathbf{y}$. Given a predictor $p:\mathcal{X}\to[0,1]$, we will use the function $12p-$ 1 which has range $[-1,1]$ when considering correlation. The notion of weak agnostic learning we use is along the lines of that defined in [34, 30, 13].
@@
returns a hypothesis $h$ such that with probability at least $0.99$, [^10] As the algorithm has access to random examples, for any $\rho>0$, we can always boost the success probability to guarantee that for any $0<\tau\leq\beta$, the output of the algorithm will satisfy, with probability at least $1-\rho$, $\mathrm{cor}_{\mathcal{D}}(\mathbf{y},h(\mathbf{x}))\geq\beta-\tau$. The increase in sample complexity and runtime bounded by a factor that is polynomial in $1/\tau,\log(1/\rho)$.
@@
  • The most powerful weak agnostic learner is one that can handle any positive $\alpha$, and has $\beta=\operatorname{poly}(\alpha)$. Colloquially weak agnostic learning is sometimes used to refer to this setting (where $\alpha$ approaches 00), but as resulting on agnostic boosting show [30, 13], this is in fact equivalent to strong agnostic learning.  [^11] We would expect the sample complexity and computational complexity to grow as an (ideally polynomial) function of $1/\beta$.
@@
We observe that weak learning is essentially equivalent to non-trivial square loss minimization, that beats random guessing. This observation is not entirely new, but will help elucidate several of our results.
@@ 3.2 Multigroup Fairness Notions
$\mathsf{ECE}(p)=\operatorname*{\mathbf{E}}\left[\Bigl{\lvert}\operatorname*{\mathbf{E}}[\mathbf{y}|p(\mathbf{x})]-p(\mathbf{x})\Bigr{\rvert}\right].$ (3)
@@
$\max_{c\in\mathcal{C}}\operatorname*{\mathbf{E}}[|c(\mathbf{x})(\operatorname*{\mathbf{E}}[\mathbf{y}|p(\mathbf{x})]-p(\mathbf{x}))|]\leq\tau.$
@@
Thus, if $\bar{\mathcal{C}}$ is the hypothesis class representing the {±$\{\pm 1\}$ versions of the functions in $\mathcal{C}$, then $p$ is $(\bar{\mathcal{C}},3\tau)$-multiaccurate with respect to $\mathcal{D}$ as per Definition 3.6. As a result, we will move freely between the $\{0,1\}$ and {±$\{\pm 1\}$ versions of functions.
@@ 4.1 Multiaccuracy and Weak Agnostic Learning
  • For any $k:[0,1]\to[-1,1]$, $k(p)$ is not a $(1/2,\beta)$ weak agnostic learner for $\mathcal{C}$ for any $\beta>0$.
@@
We consider the uniform distribution on the domain {±$\{\pm 1\}^{n}$. Let $\operatorname*{\mathbf{E}}[\mathbf{y}|\mathbf{x}]=p^{*}(\mathbf{x})$ for $p^{*}:\{\pm 1\}^{n}\to[0,1]$. Every such function specifies a joint distribution $\mathcal{D}$ on {±$\{\pm 1\}^{n}\times\{0,1\}$.
Let $\mathcal{C}=\{c:\{\pm 1\}^{n-1}\to[-1,1]\}$ be the set of all functions of the first $n-1$ bits. Now the the predictor $p(x_{1},\ldots,x_{n})=p^{*}(x_{1},\ldots,x_{n-1},-x_{n})$. If we write the multilinear expansion of $p$ as
@@
It follows that for any function $c(x_{1},\ldots,x_{n-1})$ of the first $n-1$ bits,
@@
since both $c$ and $p^{*}_{1}$ only depend on the first $n-1$ bits. So $p$ is $(\mathcal{C},0)$-multiaccurate for $\mathcal{D}$.
@@
For any pair of distinct indices $i,j\in[n-1]$ let us denote the distribution $\mathcal{D}_{ij}$ where the marginal over $\mathbf{x}$ is unfiorm and the conditional label distribution $\mathbf{y}|\mathbf{x}$ is defined by the Bayes optimal predictor
@@
For any $\mathcal{C},\mathcal{D}$, on can post-process a $(\mathcal{C},\tau)$-multiaccurate predictor to obtain a $(\alpha,2\alpha-1-2\tau)$-weak agnostic learning for $\mathcal{C}$ under the distribution $\mathcal{D}$. In particular, any $(\mathcal{C},\tau)$-multiaccurate predictor $p$ for $\mathcal{D}$ satisfies
@@
$=\operatorname*{\mathbf{E}}_{\mathcal{D}}[c(x)(p^{*}(\mathbf{x})p(\mathbf{x})-(1-p^{*}(\mathbf{x}))(1-p(\mathbf{x}))]+\operatorname*{\mathbf{E}}_{\mathcal{D}}[c(\mathbf{x})(p^{*}(\mathbf{x})-p(\mathbf{x}))]$

### v2

@@ 2.2 Hardcore Sets of Optimal Density
A measure on $\mathcal{X}$ is a function $\mu:\mathcal{X}\rightarrow[0,1]$, which is not identically 0. It defines a probability distribution $\bar{\mu}$ by re-weighting $\mathcal{D}_{\mathcal{X}}$ by $\mu$. A hardcore measure $\mu$ is one which induces a hardcore distribution $\bar{\mu}$. Given a measure $\mu:\mathcal{X}\rightarrow[0,1]$ and a base distribution $\mathcal{D}_{\mathcal{X}}$ on $\mathcal{X}$, the density of $\mu$ in $\mathcal{D}_{\mathcal{X}}$ is given by
@@
IHCL++++ from multicalibration.
Recently, Casacuberta, Dwork and Vadhan [8] showed that starting from a multicalibrated predictor, one can derive a stronger and more general version of the IHCL, which they call IHCL++++. Specifically, they show that given a multicalibrated partition, we can construct hardcore distributions of locally optimal density on each (large enough) piece of the partition. As expected, given that multicalibration is stronger than multiaccuracy, one can recover the original IHCL from IHCL++++ by “gluing” these local hardcore distributions. This glued distribution is a global $2\delta$-dense hardcore measure.
@@
We show how to construct hardcore distributions with optimal density $2\delta$ as in [8] but with the lower complexity bound of $q=O(1/(\varepsilon^{2}\delta^{2}))$ for $\mathcal{C}_{q}$ as in [46]. Indeed, we give a simple and explicit form for such a measure, starting from a predictor satisfying calibration and a weighted variant of multiaccuracy that we call weighted multiaccuracy. Both calibration and weighted multiaccuracy are implied by multicalibration, so this gives an alternate proof of the optimal density result of [8]. But crucially, the cost of obtaining both these conditions together (by many measures, including oracle calls to the weak learner) is not much more than that of multiaccuracy, which yields a considerably better bound on the parameter $q$. To formally state our result, we define weighted multiaccuracy.
We will consider weight functions $w:[0,1]\to[1,2]$, which assign weights to predictions, as is standard in the calibration literature [19, 5]. While the range [1,1][-1,1] is more standard, this choice of range is appropriate for hardcore measures. The weight function that will be important for us is
@@
Along the way, our argument fixes a gap in the gluing argument used to derive the IHCL with density $2\delta$ from IHCL++++ in [8], see the discussion at the end of Section 5.3.666Gluing is straightforward for hardcore sets, but more subtle for hardcore measures.
@@
We explain at an intuitive level how $\mu_{\mathsf{Max}}$ improves over $\mu_{\mathsf{TTV}}$ and gets optimal density. Consider a level set $S_{v}=\{x:p(x)=v\}$ of the predictor $p$ and partition it into $S_{v}^{0}$ and $S_{v}^{1}$ based on whether $g(x)$ is 0 or $1$. $\mu_{\mathsf{TTV}}$ assigns weight 1v1-v to the points in $S_{v}^{1}$ and $v$ to points in $S_{v}^{0}$. If the predictor $p$ is calibrated for this level set, so that $\Pr[g(\mathbf{x})=1|\mathbf{x}\in S_{v}]=v$, then
@@
Crucially, the measure gives equal weight v(1v)v(1-v) to $S_{v}^{0}$ and $S_{v}^{1}$.
Now consider the measure $\mu_{\mathsf{Max}}$. If $v<1/2$ then $\mu_{\mathsf{Max}}$ assigns weight $1$ to points in $S_{v}^{1}$, which are in the minority, and constitute a $v$ fraction of $S_{v}$. It assigns weight v/(1v)v/(1-v) to points in $S_{v}^{0}$, which are in the majority, and constitute a 1v1-v fraction of $S_{v}$. Overall the weight given to each of $S_{v}^{0}$ and $S_{v}^{1}$ is $v=\min(v,1-v)$. When $v>1/2$, the set $S_{v}^{1}$ has the majority of points, so $\mu_{\mathsf{Max}}$ assigns weight $1$ to points in $S_{v}^{0}$, and weighs $S_{v}^{1}$ so that both have equal weight $\min(v,1-v)$.
Why is this a good weighting strategy? Consider the majority predictor $\tilde{p}=\mathbf{1}[p\geq 1/2]$. [^7] This predictor is the best-response if $p$ is Bayes optimal. $\mu_{\mathsf{Max}}$ assigns weight $1$ to the points where $\tilde{p}$ is incorrect. The rest of the weights are assigned so that $S_{v}$ is balanced, assuming that $p$ is in fact calibrated. This clearly ensures that $\tilde{p}$ is incorrect with probability $1/2$. We now want to argue both that $\mu_{\mathsf{Max}}$ has good density and that no hypothesis in $\mathcal{C}$ can do better than $1/2$. It is intuitive that if $\mu_{\mathsf{Max}}$ is hard for $\tilde{p}$, which incorporates the distinguishing power of the class $\mathcal{C}$, then it should also be hard for $\mathcal{C}$.
@@
It remains to argue that the measure $\mu_{\mathsf{Max}}$ is hardcore. In Lemma 5.8 we present a characterization of the prediction advantage of $\mathcal{C}$ under the distribution $\bar{\mu}_{w}$ in terms of the $(w,\mathcal{C})$-weighted multiaccuracy error of $p$ under the distribution $\mathcal{D}^{g}$. This tells us that if the multiaccuracy error is $\tau$, then the prediction advantage under $\bar{\mu}_{w}$ is $\tau/\mathsf{dns}(\mu_{w})$. This proof technique is different from [8, Lemma 2.17] who show that if a function is indistinguishable from a constant function $v$ to $\mathcal{C}$ on each level set $S_{v}$, then then one can obtain a hardcore set of density $2\min(v,1-v)$ within $S_{v}$. The required indistinguishability holds for every level set by the assumption of multicalibration. Our result can be seen as a simplification and extension of the hardness result in [46] to weighted measures.
Note that calibrated multiaccuracy does not give the stronger IHCL++++ theorem proved in [8]; indeed that seems to require the full power of multicalibration. Also, the min-max based proof of Nisan as cited in [28] and the boosting-based proof of [35] give $q$ which grows as $O(\log(1/\delta))$, but they do not get optimal density. The $1/\varepsilon^{2}$ dependence on $\varepsilon$ was recently showed to be necessary by [4].
@@ 2.3 Multiaccuracy and Restricted Weak Agnostic Learning
  • The Blue line with slope 11 follows from the boosting results of [30, 13].
@@ 3.1 Agnostic Learning
The transformation $\mathbf{y}\to 2\mathbf{y}-1$ maps $\{0,1\}$ to $\{\pm 1\}$. Hence the correlation measures the correlation between $c$ and $\{\pm 1\}$ labels associated with $\mathbf{y}$. Given a predictor $p:\mathcal{X}\to[0,1]$, we will use the function 2p12p-1 which has range [1,1][-1,1] when considering correlation. The notion of weak agnostic learning we use is along the lines of that defined in [31, 30, 13].
@@
returns a hypothesis $h$ such that with probability at least 0.990.99, [^10] As the algorithm has access to random examples, for any $\rho>0$, we can always boost the success probability to guarantee that for any $0<\tau\leq\beta$, the output of the algorithm will satisfy, with probability at least $1-\rho$, $\mathrm{cor}_{\mathcal{D}}(\mathbf{y},h(\mathbf{x}))\geq\beta-\tau$. The increase in sample complexity and runtime bounded by a factor that is polynomial in $1/\tau,\log(1/\rho)$.
@@
  • The most powerful weak agnostic learner is one that can handle any positive $\alpha$, and has $\beta=\operatorname{poly}(\alpha)$. Colloquially weak agnostic learning is sometimes used to refer to this setting (where $\alpha$ approaches 0), but as results on agnostic boosting show [30, 13], this is in fact equivalent to strong agnostic learning. [^11] We would expect the sample complexity and computational complexity to grow as an (ideally polynomial) function of $1/\beta$.
@@
We observe that weak learning is essentially equivalent to non-trivial squared loss minimization that beats random guessing. This observation is not entirely new, but will help elucidate several of our results.
@@ 3.2 Multigroup Fairness Notions
[…]
@@
[…]
@@
Thus, if $\bar{\mathcal{C}}$ is the hypothesis class representing the $\{\pm 1\}$ versions of the functions in $\mathcal{C}$, then $p$ is $(\bar{\mathcal{C}},3\tau)$-multiaccurate with respect to $\mathcal{D}$ as per Definition 3.6. As a result, we will move freely between the $\{0,1\}$ and $\{\pm 1\}$ versions of functions.
@@ 4.1 Multiaccuracy and Weak Agnostic Learning
  • For any $k:[0,1]\to[-1,1]$, $k(p)$ is not a $(1/2,\beta)$-weak agnostic learner for $\mathcal{C}$ for any $\beta>0$.
@@
We consider the uniform distribution on the domain $\{\pm 1\}^{n}$. Let $\operatorname*{\mathbf{E}}[\mathbf{y}|\mathbf{x}]=p^{*}(\mathbf{x})$ for $p^{*}:\{\pm 1\}^{n}\to[0,1]$. Every such function specifies a joint distribution $\mathcal{D}$ on $\{\pm 1\}^{n}\times\{0,1\}$.
Let $\mathcal{C}=\{c:\{\pm 1\}^{n-1}\to[-1,1]\}$ be the set of all functions of the first n1n-1 bits. Now define the predictor $p$ as $p(x_{1},\ldots,x_{n})=p^{*}(x_{1},\ldots,x_{n-1},-x_{n})$. If we write the multilinear expansion of $p^{*}$ as
@@
It follows that for any function $c(x_{1},\ldots,x_{n-1})$ of the first n1n-1 bits,
@@
since both $c$ and $p^{*}_{1}$ only depend on the first n1n-1 bits. So $p$ is $(\mathcal{C},0)$-multiaccurate for $\mathcal{D}$.
@@
For any pair of distinct indices $i,j\in[n-1]$ let us denote the distribution $\mathcal{D}_{ij}$ where the marginal over $\mathbf{x}$ is uniform and the conditional label distribution $\mathbf{y}|\mathbf{x}$ is defined by the Bayes optimal predictor
@@
Our next result shows that assuming the existence of pseudorandom functions that fool the class $\mathcal{C}$, a very general result can be established — namely that, in general, no post-processing of a multiaccurate predictor yields a weak agnostic learner.
Theorem 4.2.
Let $\mathcal{C}$ be a hypothesis class of functions with range $\{-1,1\}$ and $\mathcal{D}_{\mathcal{X}}$ be a distribution over $\mathcal{X}$. Suppose there exists an efficiently computable function $g:\mathcal{X}\rightarrow\{0,1\}$ such that,
[…]
Then there exists a distribution $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$ with the marginal distribution over $\mathcal{X}$ being $\mathcal{D}_{\mathcal{X}}$, and an efficiently computable predictor $p:\mathcal{X}\rightarrow[0,1]$ such that,
  • $p$ is $(\mathcal{C},\tau)$-multiaccurate for $\mathcal{D}$.
  • For any $k:[0,1]\rightarrow[-1,1]$, $k(p)$ is not a $(q-\tau,\beta)$ agnostic learner for $\mathcal{C}$ for any $\beta>\tau$, where
    $q=\max_{c\in C}\left\{\min\{\Pr_{\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})=1],\Pr_{\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})=-1]\}\right\}.$
Proof.
Fix some $f\in\mathcal{C}$ and let $q=\Pr_{\mathcal{D}_{\mathcal{X}}}[f(\mathbf{x})=1]$. Assume without loss of generality that $0<q\leq\frac{1}{2}$. Let $\widehat{f}:\mathcal{X}\rightarrow\{-1,1\}$ be a (randomized) function such that
[…]
We have defined $\widehat{f}$ to be a $\pm 1$-valued function that is a balanced version of $f$; in particular, $\Pr_{\mathcal{D}_{\mathcal{X}}}[\widehat{f}(\mathbf{x})=1]=\frac{1}{2}$. Note that
$\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[f(\mathbf{x})\widehat{f}(\mathbf{x})]$ $=\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[f(\mathbf{x})\widehat{f}(\mathbf{x})|f(\mathbf{x})=1]\cdot q+\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[f(\mathbf{x})\widehat{f}(\mathbf{x})|f(\mathbf{x})=-1]\cdot(1-q)=2q.$ (6)
Let $\widehat{g}:\mathcal{X}\rightarrow\{-1,1\}$ be defined as the $\pm 1$-valued version of $g$, i.e. $\widehat{g}(x)=2g(x)-1$. Define the distribution $\mathcal{D}$ with the marginal distribution $\mathcal{D}_{\mathcal{X}}$ over $\mathcal{X}$ such that,
$\operatorname*{\mathbf{E}}[\mathbf{y}|\mathbf{x}]=p^{*}(\mathbf{x})=\frac{2+\widehat{f}(\mathbf{x})+\widehat{g}(\mathbf{x})}{4}.$
Finally, define a predictor $p:\mathcal{X}\rightarrow[0,1]$ as
$p(x)=\frac{2+\widehat{f}(x)-\widehat{g}(x)}{4}.$
In order to check that $p$ is $(\mathcal{C},\tau)$-multiaccurate for $\mathcal{D}$, we can see that for any $c\in C$,
$\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})(p^{*}(\mathbf{x})-p(\mathbf{x}))]$ $=\frac{1}{2}\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})\widehat{g}(\mathbf{x})]\leq\tau.$
The latter inequality follows from the condition on $g$ in the statement of the theorem.
Similary, we can check that,
$\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})(p^{*}(\mathbf{x})-p(\mathbf{x}))]$ $=\frac{1}{2}\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[c(\mathbf{x})\widehat{g}(\mathbf{x})]\geq-\tau.$
Thus, $p$ is $(C,\tau)$-multiaccurate.
We observe that $p(x)\in\left\{0,\frac{1}{2},1\right\}$. For any such $v$ in the range of $p$, as we have picked $f\in C$, the condition on $g$ in the theorem statement shows that $|\mathrm{cor}(\mathbf{y},p(\mathbf{x})|p(\mathbf{x})=v)|\leq\tau$.
On the other hand, we can also observe that (using Eqn. (6)),
$\mathrm{cor}(\mathbf{y},f)$ $=\operatorname*{\mathbf{E}}[(2\mathbf{y}-1)f(\mathbf{x})]$
$=\operatorname*{\mathbf{E}}[(2p^{*}(\mathbf{x})-1)f(\mathbf{x})]$
$=\frac{1}{2}\operatorname*{\mathbf{E}}[(\widehat{f}(\mathbf{x})+\widehat{g}(\mathbf{x}))f(\mathbf{x})]$
$\geq q-\tau.$
Thus as long as $q>\tau$, there is a function in $\mathcal{C}$ with non-trivial correlation with $\mathbf{y}$, but no post-processing of $p$ yields a weak agnostic learner with correlation better than $\tau$. (Note that if $q$ is almost 0 for every $f\in\mathcal{C}$, then $\mathcal{C}$ only contains almost constant (w.r.t $\mathcal{D}_{\mathcal{X}}$) functions.) ∎
Remark 4.3.
The existence of a $g$ satisfying the conditions in the statement of Theorem 4.2 with $\tau$ being negligible follows from the existence of one-way functions as long as $\mathcal{C}$ is contained in the class of polynomial-sized circuits [15]. We have assumed above that $p$ can be a randomized predictor; again this assumption can be removed assuming the existence of one-way functions. Also note that if every $c\in\mathcal{C}$ is balanced, then we get $q=1/2$; this was the case with the class used in Theorem 4.1.
@@
For any $\mathcal{C},\mathcal{D}$, one can post-process a $(\mathcal{C},\tau)$-multiaccurate predictor to obtain a $(\alpha,2\alpha-1-2\tau)$-weak agnostic learning for $\mathcal{C}$ under the distribution $\mathcal{D}$. In particular, any $(\mathcal{C},\tau)$-multiaccurate predictor $p$ for $\mathcal{D}$ satisfies
@@
$=\operatorname*{\mathbf{E}}_{\mathcal{D}}[c(\mathbf{x})(p^{*}(\mathbf{x})p(\mathbf{x})-(1-p^{*}(\mathbf{x}))(1-p(\mathbf{x}))]+\operatorname*{\mathbf{E}}_{\mathcal{D}}[c(\mathbf{x})(p^{*}(\mathbf{x})-p(\mathbf{x}))]$